how to calculate root mean square of bromine molecule at 23°c and 1 atm
1
Expert's answer
2019-04-30T04:30:37-0400
"v_{rms}=\sqrt{\smash[b]{\frac{3RT}{M}}}"

Where

"T(K)= 23^\circ C +273.15= 296.15 K"

"M(Br_2)= 79.9\times2= 159.8 \frac{g}{mol}"

"v_{rms}=\sqrt{\smash[b]{\frac{3\times 8.314\times 296.15}{159.8\times10^-3}}} = 215 \frac{m}{s}"